OpenStudy (anonymous):

1. What is the probability that a card drawn at random from a shuffled deck of 52 normal playing cards is a Heart or a Faced card? Be sure to avoid double counting. 2. how many arrangements can be made from the 26 letters in the English alphabet by: a. 2 different letters? b. 3 different letters? c. 4 letters without replacement? d. 4 letters with replacement? 3. Fourteen people try out for a baseball team. In how many different ways could they select: a. the pitcher and the catcher? b. the three outfielder, after the pitcher and catcher have been selected?

OpenStudy (anonymous):

3 c. the four infielders (1st, 2nd, SS, and 3rd), after the other five team members have been selected?

OpenStudy (anonymous):

do you know how many hearts are in the deck?

OpenStudy (anonymous):

@satellite73 no

OpenStudy (venomblast):

there's 13 hearts 13 spades, 13 diamonds and 13 Clubs

OpenStudy (venomblast):

For faces, there's 4 kings, 4 queens and 4 Jacks. P(FACE) = 12/52

OpenStudy (venomblast):

THis is how you set it up. P(FUH) F= Face and H= heart P(FUH)=P(F)+P(H)-P(FH) 12/52 + 13/52 - 1/52 = 24/52 = 6/13

OpenStudy (venomblast):

@mgmk Please open another question.